OpenStudy (anonymous):

5-2t+sin(8t)/5t+cos(8t) find the limit as t goes to negative infinity

OpenStudy (anonymous):

is that \[\frac{5-2t+\sin(8t)}{5t+\cos(8t)}\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

don't mean to rush but i have 6 min to put in an answer on my online hw before it's considered late LOL procrastination at it's best

OpenStudy (anonymous):

-2/5

OpenStudy (anonymous):

OMG I LOVE YOU SO MUCH RIGHT NOW

OpenStudy (zzr0ck3r):

divide everything by t, you will see that after you run the limit, by the squeeze theorem that the trig functions over t will go to 0 and the constants will go to 0

OpenStudy (anonymous):

that was close ok now how did you come to that lol

OpenStudy (zzr0ck3r):

he went to wolfram alpha

OpenStudy (anonymous):

Yes :3

OpenStudy (anonymous):

so sin8t/t is 0

OpenStudy (anonymous):

same with cos8t/t

OpenStudy (zzr0ck3r):

\(\frac{5-2t+sin(8t)}{5t+cos(8t)} =\frac{\frac{5}{t}-2+\frac{sin(8t)}{t}}{5+\frac{cos(8t)}{t}}\)

OpenStudy (zzr0ck3r):

take the limit and every term with a t under it goes to 0

OpenStudy (zzr0ck3r):

you are left with -2/5

OpenStudy (anonymous):

is that because the denominator is going infinitly big?

OpenStudy (zzr0ck3r):

\(-1\le \cos(ax) \le 1 \implies \frac{-1}{t}\le\frac{cos(ax)}{t}\le\frac{1}{t} \) taking the limit we see \(0\le\frac{cos(ax)}{t}\le0\) thus it must be 0